2) To analyze the chromium in an unknown samp 5623 e sample was pretreated and diluted to the exact 100 m ark in a volumetric flask. Two 20.0 mD samples of this solution were taken. In the first sample, a Cr-EDTA complex is developed and sufficient water is added to give a total volume 00.0 mL. In the s complex. water is concentration of 15 m as added before developing the Cr-EDTA Sufficient c also added to give a total of absorbance of each is measured in a 1.0 cuvette at 450 nm and is found to be for the first solution and 0.30for the second solution. Calculate the chromium (atomic wt 52.0 g/mol) in the original sample. (10 ram chro mium s chromium S62? 3) Molar absorptivity of cobalt and copper complexes with a ligand Lane Eco- 605 and Ecy 552 at 510 nm and goo 750 625 o g sample containing Co and Cu was dissolved and dilute to 500.0 m A 25.0-mL aliquot was treated and complexed with the ligand L. This solution had an of 0.646 at at 625 cell. Calculate the percentage of Co (Mw: and cu Mw: 6 g/mo) in the original 58.9 g/mol) (11 At Sion m o 646 60 Ca 5520 cu At 6 RSAm 646 -SS 2 Cau 605 o 462
Answer #1

2) Analysis of chromium

20 ml aliquot diluted to 100 ml without standard gave aborbance = 0.15

IInd 20 ml aliquot with standard,

concentration of standard in diluted sample = 15 mM x 10 ml/100 ml = 1.5 mM

let x be the concentration of Cr in diluted sample

then,

x/(x + 1.5) = 0.15/0.30

0.30x = 0.15x + 0.225

x = 0.225/0.15 = 1.5 mM

concentration of Cr in original sample = 1.5 x 100/20 = 7.5 mM

grams of Cr = 7.5 x 52 x 0.1/1000 = 0.039 g

%Cr in original sample = 0.039 x 100/1.5623 = 2.50%

3) Absorbance of mixture is sum total of absorbances of all the species in solution

Absorbance = molar absorptivity x path length x concentration

let A be concentration of Co and B be concentration of Cu

then,

at 510 nm,

0.646 = 605[A] + 552[B]

0.462 = 121[A] + 750[B]

Solving,

[A] = 5.93 x 10^-4 M

[B] = 5.20 x 10^-4 M

%Co in sample = 5.93 x 10^-4 x 0.5 x 58.9 x 100/0.425 = 4.11%

%Cu in sample = 5.20 x 10^-4 x 0.5 x 63.5 x 100/0.425 = 3.88%